solving quadratic equations by factoring
y=x squared+6x+8
5wsquared=30w

Answer by stanbon(75887)

solving quadratic equations by factoring
y=x squared+6x+8
x^2 +6x + 8 = 0
(x+4)(x+2) = 0
x = -4 or x = -2
----------------------------
5wsquared=30w
5w^2 - 30w = 0
5w(w-6) = 0
w = 0 or w = 6
---------------------
Cheers,
Stan H.
---------------------